OpenStudy (anonymous):

Solve the quadratic equation by factoring. 21x^2 + 3x = 0

OpenStudy (ybarrap):

3(7x^2 + x) = 0 7x^2 + x = 0 7x(x + 1/7) = 0 Ans x = 0 and x = -1/7